OpenStudy (anonymous):

Find the first four terms of the geometric sequence with t1 = 16 and r = 1/2

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

@Eyad

OpenStudy (anonymous):

T1=16 , r=1/2 Hint: ==== T1=A=First term . Apply the general formula : Tn=ar^n-1 and get the first foir terms :)

OpenStudy (anonymous):

four *

OpenStudy (anonymous):

16^1/2-1?

OpenStudy (anonymous):

n=?

OpenStudy (anonymous):

n=4

OpenStudy (anonymous):

n is the number of the term Like : T1=ar^1-1 T2=ar^2-1 T3=ar^3-1 ...

OpenStudy (anonymous):

Got it ?

OpenStudy (anonymous):

16(1/2)^4-1

OpenStudy (anonymous):

NO I dont got how to get the next for terms?

OpenStudy (anonymous):

four*

OpenStudy (anonymous):

t1=a=16 T2=ar^n-1=16*1/2^2-1=? T3=ar^n-1=16*1/2^3-1= ? Got it ?

OpenStudy (anonymous):

t1=16,t2=3 t3=1

OpenStudy (anonymous):

t4=-5/8

OpenStudy (anonymous):

Cool :)

OpenStudy (anonymous):

:)